In exponential form of a complex number how ,e raised to power iota sin theta = cos (sin theta)+ iota sin (sin theta)?

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
e^ix = cos x + i sinx
Now
 
e ^i sinx = cos(sin x) + i sin(sin x)
 
Hence proved
 
Regards
Arun (askIITians forum expert)